OpenStudy (anonymous):

what is the uniit rate 324 words in 6 minutes, round to the nearest hundreth if necessary

OpenStudy (shaik0124):

The unit rate of 324 words in 6 minutes rounded to the nearest hundredth would be 54.72

jimthompson5910 (jim_thompson5910):

You simply divide the number of words by the number of minutes.A better example is probably 100 words in 2 minutes = 100/2 = 50 words per minute

OpenStudy (shaik0124):

324/6 =54 words per mintue